OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 7x2 + 1 = 5x double root real and rational root real and irrational root imaginary root

OpenStudy (mathmale):

Hi, @magan , Your best bet is to actually find the roots of 7x2 + 1 = 5x, or\[7x^2 + 1 = 5x\] First, put the equation into standard form by moving that 5x term to the left side. To do this, subtract 5x from both sides. Next, compare the resulting equation to \[ax^2 + bx + c = 0\] and determine what the values of a, b and c are. What are they? a= ? b= ? c= ?